OpenStudy (anonymous):

Find the values of x that give relative extrema for the function f(x) = 3x^5 - 5x^3.

OpenStudy (turingtest):

can you take the derivative of this function?

OpenStudy (anonymous):

yes its 15x^4 - 15x^2

OpenStudy (turingtest):

great solving for zero will give you some of the critical numbers have you got that yet?

OpenStudy (anonymous):

no but i think i did something wrong because as i equaled it to 0 i got x^3= x

OpenStudy (turingtest):

well zero is one answer, but the best way to do it is to factor out 15x^2 giving 15x^2(x^2-1)=0 then the roots should be obvious

OpenStudy (anonymous):

so is the relative maximum x=0 and relative minima x= plus or minus 1?

OpenStudy (turingtest):

well we haven't checked which is which at least I haven't, hold on...

OpenStudy (turingtest):

the maximum occurs at x=-1 what made you think it was at zero?

OpenStudy (turingtest):

how did you check which was a maxima and which was a minima?

OpenStudy (anonymous):

i made a chart but i think i did it wrong

OpenStudy (turingtest):

Breaking the solutions up into intervals and testing the value of the derivative in each interval can tell you which is which, is that what you meant by a chart?

OpenStudy (anonymous):

yes exactly that

OpenStudy (turingtest):

look at the space on the side of x<-1 take any point like x=-2 and plug it into the derivative f'(x) if f'(-2)>0 then f(x) is increasing on that interval if f'(-2)<0 then f(x) is decreasing on that interval if the function changes from increasing to decreasing then it is a maximum what did you get for that test?

OpenStudy (anonymous):

mkay hold on my work is all over the place

OpenStudy (turingtest):

if you can't find your work just check now plug x=-2 into f(x)=15x^4-15x^2

OpenStudy (anonymous):

ok thank you so much for your help honestly i appreciate it because i have a calc packet to work on and i'm really quite clueless on more than half of it

OpenStudy (turingtest):

anytime let me know if you have any questions.

OpenStudy (anonymous):

ok i got 809,100

OpenStudy (turingtest):

lol well the important thing is that it is >0 so we have the fact that f(x) is INCREASING on x<-1 so that is one interval taken care of what did you get for a number in the next interval? (-1,0)

OpenStudy (turingtest):

and you should have gotten 180 for x=-2... f'(-2)=15(-2)^4-15(-2)^2=240-60=180

OpenStudy (turingtest):

Now that we know the first interval is decreasing pick any point from the next one (-1,0) say -1/2 or -1/4, something like that

OpenStudy (anonymous):

ok i did 15x-2 and then placed the answer to the fourth power that's why

OpenStudy (turingtest):

I see, it happens so what about the next interval pick a value for x and check f'(x) in the interval (-1,0)

OpenStudy (anonymous):

ok i got -2.81

OpenStudy (turingtest):

sounds reasonable and what does that tell us about f(x) in that interval?

OpenStudy (anonymous):

its decreasing

OpenStudy (turingtest):

right, so we can make a chart...

OpenStudy (turingtest):

so it looks like f(x) changed from increasing to decreasing at x=-1so what does that tell us about the point f(-1) ?

OpenStudy (anonymous):

its the maximum

OpenStudy (turingtest):

Good! now test the next interval (0,1)

OpenStudy (turingtest):

by the way, it's "A maximum" not necessarily the only one, we haven't tested yet to see if it is local or global.

OpenStudy (anonymous):

ohh ok between 0 and 1 i got a negative

OpenStudy (turingtest):

right, so we could see in our chart no change at x=0so we learn nothing. What about the last interval?